My answer(s) are incorrect. Please help. The Ka values for H2CO3 are: Ka1=4.45x10^-7 and Ka2=4.69x10^-11.How would you prepare 1.20 L of a buffer with a pH of 9.13 from 0.270 M Na2CO3 and 0.190 M HCl? (This problem requires values
Answer #1

The reaction between Na2CO3 and HCl: Na2CO3 + HCl ----> NaHCO3 + NaCl

pKa2 of NaCO3 = -Log(Ka2) = -Log(4.69x10-11) = 10.33

According to the Henderson-Hasselbulch equation:

pH = pKa + Log([Na2CO3]/[NaHCO3])

i.e. 9.13 = 10.33 + Log([Na2CO3]/[NaHCO3])

i.e. Log([Na2CO3]/[NaHCO3]) = -1.2

i.e. [Na2CO3]/[NaHCO3] = 10-1.2 = 0.0631

Now, the pecent fraction of Na2CO3 = {0.0631/(1+0.0631)}*100 = 5.9%

And the percent fraction of NaHCO3 = 100-5.9 = 94.1%

Accordingly, you can add HCl to the Na2CO3 solution.